1)Select the approximate values of x that are solutions to f(x) = 0, where
f(x) = -9x2 + 4x + 7

2)Select the approximate values of x that are solutions to f(x) = 0, where
f(x) = -2x2 + 3x + 9


Answer by stanbon(75887)

1)Select the approximate values of x that are solutions to f(x) = 0, where
f(x) = -9x2 + 4x + 7
-------
-9x^2 + 4x + 7 = 0
Use the Quadratic formula to get:
x = [-4 +- sqrt(16-4*-9*7))/(2*-9)
---
x = [-4 +- sqrt(268))/(-18)
---
x = [-4 +- 2sqrt(67))/-18
----
x [-2 +- sqrt(67))(-9)
==================================
2)Select the approximate values of x that are solutions to f(x) = 0, where
f(x) = -2x2 + 3x + 9
-2x^2 + 3x + 9 = 0
-2x^2 -6x+3x + 9 = 0
-2x(x+3)+3(x+3) = 0
(x+3)(-2x+3) = 0
x = -3 or x = 3/2
--------------------------
Cheers,
Stan H.
=====================